# Approximating the universal thermal climate index using sparse regression with orthogonal polynomials

### v1

@@ Abstract
This article explores novel data-driven modeling approaches for analyzing and approximating the Universal Thermal Climate Index (UTCI), a physiologically-based metric integrating multiple atmospheric variables to assess thermal comfort. Given the nonlinear, multivariate structure of UTCI, we investigate symbolic and sparse regression techniques as tools for interpretable and efficient function approximation. In particular, we highlight the benefits of using orthogonal polynomial bases-such as Legendre polynomials-in sparse regression frameworks, demonstrating their advantages in stability, convergence, and hierarchical interpretability compared to standard polynomial expansions. We demonstrate that our models achieve significantly lower root-mean squared losses than the widely used sixth-degree polynomial benchmark-while using the same or fewer parameters. By leveraging Legendre polynomial bases, we construct models that efficiently populate a Pareto front of accuracy versus complexity and exhibit stable, hierarchical coefficient structures across varying model capacities. Training on just 20% of the data, our models generalize robustly to the remaining 80%, with consistent performance under bootstrapping. The decomposition effectively approximates the UTCI as a Fourier-like expansion in an orthogonal basis, yielding results near the theoretical optimum in the $L_{2}$ (least squares) sense. We also connect these findings to the broader context of equation discovery in environmental modeling, referencing probabilistic grammar-based methods that enforce domain consistency and compactness in symbolic expressions. Taken together, these results illustrate how combining sparsity, orthogonality, and symbolic structure enables robust, interpretable modeling of complex environmental indices like UTCI - and significantly outperforms the state-of-the-art approximation in both accuracy and efficiency.
Keywords:
Climate Index Sparse regression Orthogonal polynomials
@@ 1 Introduction
The increasing availability of high-resolution environmental data—driven by advances in remote sensing, sensor networks, and global monitoring—has created new opportunities and challenges for climate analysis [26, 45, 46, 47, 48, 49, 50]. These complex, spatiotemporal datasets offer insights into atmospheric, hydrological, and ecological processes [28, 35, 41, 51, 53], but often exceed the capacity of traditional methods. Machine learning (ML) provides tools well-suited for extracting structure from such data [22, 59, 64], with applications ranging from forecasting to anomaly detection [33, 34]. One emerging use of ML is in refining climate indices—simplified metrics like the Universal Thermal Climate Index (UTCI), Palmer Drought Severity Index (PDSI), or Oceanic Niño Index (ONI)-that summarize complex climate phenomena [24]. The UTCI is a comprehensive bioclimatic index developed to assess the thermal environment with direct relevance to human thermal physiology. The UTCI is expressed in degrees Celsius and is intended to be interpreted analogously to air temperature, though with a more nuanced and physiologically grounded foundation.
The UTCI uses a detailed thermo-physiological model, see Fig. 1, to simulate human thermal responses—like sweating, shivering, and skin blood flow-under changing conditions. This enables more accurate assessments than traditional indices. UTCI categorizes thermal stress on a continuous scale from “extreme cold” to “extreme heat,” aiding public health communication and risk management. Widely used in biometeorology, urban planning, and occupational safety, it supports studies on urban heat islands, outdoor labor, and heat-health warnings. Its standardized approach allows international comparisons and long-term climate analysis, informing strategies for thermal comfort, public safety, and climate adaptation. It is also increasingly used in many countries, especially in Europe, as a measure of thermal comfort for outdoor conditions and is calculated as part of the operational meteorological forecast [15].
Figure 1: Schematic of the UTCI framework [31]: meteorological inputs and a clothing model feed into a physiological model to compute thermal stress. The resulting UTCI value is classified on a color scale from extreme cold to extreme heat stress.
The dataset [11] we use provides tabulated UTCI values computed using the full thermo-physiological model, covering a wide range of environmental variables: air temperature, wind, incoming radiation, and humidity. Formally, the UTCI is defined as: UTCI = Ta + Offset, where Ta is the air temperature and the Offset is the physiologically equivalent temperature difference, representing how other environmental factors modify the effect of the thermal stress on the human body. To approximate this data, the authors also introduced a sixth-degree polynomial regression model, which has since become a standard in human biometeorology due to its simplicity and computational efficiency. The goal of this study is to develop a more accurate and robust model for the UTCI Offset-one that reduces average and extreme errors, enhances interpretability, and offers fast computation for large-scale or real-time applications. Specifically, we explore the use of sparse and symbolic machine learning to approximate UTCI from meteorological inputs. Using orthogonal polynomial bases (e.g., Legendre polynomials), we develop interpretable, numerically stable models that generalize well and offer structural advantages over standard polynomial regressions.
@@
The UTCI Offset accounts for the effects of incoming radiation (expressed via mean radian temperature Tr), of wind speed (va), which influences convective heat loss, and relative humidity (rH), that affects evaporative heat loss (e.g., sweating efficiency). Furthermore, the Offset depends on Ta indirectly (since the effects of wind and humidity depend on it). In Fig. 2(a) we see how the UTCI Offset varies along the different environmental variables. Instead of the humidity (rH), the water vapor pressure (pa) can be used which is a nonlinear function of rH and the air temperature (Ta). However, the variables have different distribution, see Fig. 2(b), which impacts the extent that approximations of UTCI can generalize, discussed below.
@@
Figure 2: (a) 3D plot of UTCI Offset [11] at 5% relative humidity, showing how wind speed (va), air temperature (Ta), and mean radiant temperature difference (Tr–Ta) combine to influence thermal stress. Color indicates the UTCI Offset magnitude across these environmental dimensions. (b) The different distributions of the water vapor pressure and relative humidity in the computed Offset dataset [11]. The water vapor pressure is strongly peaked at zero, while the relative humidity is uniform across its range.
2 Methods
Equation discovery aims to learn interpretable mathematical expressions, either differential or algebraic equations, that describe dynamical systems [61, 67, 66]. Positioned at the intersection of symbolic machine learning and system identification, it is increasingly relevant in environmental and climate science, where data-driven yet transparent models are essential. Traditional models rely on expert-derived formulations [36, 37, 38, 39, 40], but the complexity and volume of climate data call for automated alternatives. Symbolic regression, which iteratively combines mathematical operators and variables to fit data, forms the core of equation discovery [60, 9, 63, 13, 16]. These methods are commonly implemented using evolutionary or other (e.g., enumerative) search algorithms [55, 54, 27, 44, 57, 6, 30].
Recent advances integrate probabilistic grammars to incorporate prior knowledge and constrain the search to physically meaningful expressions [8, 21, 7, 29, 32]. This structured approach improves both model interpretability and search efficiency, especially in domains governed by established scientific principles. Equation discovery has been applied to environmental systems [1, 2, 3, 4, 5], including ecosystem dynamics [17, 18, 19, 20, 23, 14, 42, 43, 56, 58], where it can match or surpass expert models while revealing new relationships [62, 65, 68]. Its ability to generate compact, interpretable models makes it especially suitable for climate applications, where model transparency and physical plausibility are vital.
To simplify UTCI computation, the standard approach in the literature is to use a sixth-degree regression polynomial [11, 25]. However, this method suffers from generalization issues [10]. Fig. 3(a) shows the approximation error at 5% relative humidity, while 3(b) displays a histogram of the errors, revealing a normal distribution centered at zero, indicating minimal bias. We aim to improve upon this standard approximation using equation discovery and sparse regression methods. The standard approximation uses the Offset dataset [11] and we employ the same data in our computations, see Fig. 2(a).
@@ 2 Methods
Figure 3: (a) The difference between the standard UTCI approximation and the real values for relative humidity of 5%. (b) Histogram of the differences (or errors) showing a normal distribution centered at zero.
Sparse machine learning models aim to construct parsimonious predictive functions by enforcing zero-valued coefficients in high-dimensional parameter spaces, thereby performing implicit feature selection [12]. This sparsity promotes interpretability, reduces overfitting, and improves computational tractability, especially when the number of candidate predictors is large or when strong correlations exist among inputs. Sparse regression, a key instantiation of this paradigm, extends linear regression with an $L_{1}$-norm regularization term—most notably in the Lasso (Least Absolute Shrinkage and Selection Operator)—to penalize unnecessary parameters and induce a compact representation.
@@ 3 Results and discussion
Table 1 presents a detailed comparison of model performance across a range of polynomial degrees for both standard (non-sparse) linear regression and sparse regression techniques, evaluated in the context of approximating the UTCI. The standard approximation is a sixth-degree regression polynomial model with four variables, consisting of 210 terms and achieving a root mean squared (RMS) loss of 1.12. This serves as the benchmark to be matched or improved upon. It is important to note that the standard approximation [11] does not directly employ the relative humidity (rH), but the water vapor pressure (pa) which is a highly nonlinear function derived from the relative humidity (rH) and air temperature (Ta). As we noted above, the relative humidity is well represented across its entire range, see Fig. 2(b), while the water vapor pressure is strongly peaked around zero. Optimization employing the water vapor pressure (pa) as an independent variable (instead of rH) is thus poorly conditioned and leads to instability in the regression coefficients, both in simple and sparse regression. While using the pa (instead of rH) can achieve better accuracy (lower loss), it comes at the price of losing parameter consistency across optimizations with different polynomial degrees. For this reason we report our results employing the relative humidity (rH) instead of the water vapor pressure (pa).
@@
Table 1: Root mean squared train loss, test loss and the number of parameters in approximating the UTCI Offset. The baseline reference, labeled as “Standard,” corresponds to the sixth-degree regression polynomial model with four variables. Unless otherwise stated the test loss equals the train loss. Where two loss values are reported, they indicate a notable train-test discrepancy, typically suggesting overfitting. Training is done with 20% of the data and testing is performed with 80%. Results are robust under bootstrapping.
@@
The regression methods are applied to polynomial basis expansions of increasing degree, evaluated on the basis of root mean squared (RMS) test loss and number of active parameters. Unlike many studies in the literature where models are trained on the majority of the data and evaluated on a relatively small test set, our approach inverts this paradigm: training is conducted on only 20% of the available data, while performance is assessed on the remaining 80%. Despite this stringent evaluation setting, the models achieve comparable performance on both training and test sets, underscoring their strong generalization capabilities. This performance stability is further validated through bootstrapping, which reveals minimal variance in both loss metrics and selected features across resampled datasets. The reported performance metrics—such as train/test loss and number of parameters—remain stable when the model training and evaluation process is repeated on multiple random re-samplings (bootstrapped subsets) of the data. This suggests that the results are not sensitive to specific data splits and that the models generalize well across different subsets of the dataset, indicating reliability and consistency in the reported findings. These findings demonstrate the robustness and reliability of the proposed framework.
@@
Linear regression without any sparsity constraints shows improved performance at higher degrees, with test loss reducing as model capacity increases. However, this comes with a dramatic increase in the number of parameters; it reaches over 1800 coefficients by degree 12. Furthermore, the discrepancy between train and test losses at higher degrees (e.g., 0.62 vs. 0.54 at degree 12) indicates overfitting, despite the improved predictive accuracy. The resulting models are also substantially more complex, raising concerns regarding interpretation and generalization. Sparse regression with standard polynomial bases show similar performance at low degrees but fails to converge beyond the 6th degree. This indicates that enforcing sparsity in a poorly conditioned basis becomes increasingly difficult as model complexity grows.
@@
Figure 4: Loss versus number of parameters for different polynomials degrees. The regularization parameter was varied in the lasso regression to yield a Pareto front in model accuracy and complexity for each degree.
@@
In contrast, sparse regression using an orthogonal Legendre basis exhibits superior stability and accuracy across all degrees. It outperforms the baseline 6th-degree polynomial fit from degree 8th onward, achieving a test loss of 0.88 at degree 10 with only 210 parameters—the same count as the original benchmark model, but with improved generalization. As the degree increases to 16, the loss reduces further to 0.60 using 424 parameters—a fraction of those used by the corresponding standard regression model. The orthogonality of the Legendre basis likely contributes to better numerical conditioning, facilitating sparse model discovery even at high degrees. These results emphasize the importance of basis selection and regularization strategy in symbolic regression tasks. Sparse methods, when combined with well-structured bases like Legendre polynomials, offer a promising path toward accurate, compact, and interpretable models in high-dimensional settings.
@@
Figure 5: Parameters (or polynomial coefficients) and how they change for different polynomial degrees for (a) simple regression and (c) sparse regression (using Legendre basis).
Fig. 4 illustrates the relationship between model complexity (measured by the number of parameters) and prediction accuracy (log-scaled loss) for sparse regression models using Legendre polynomial bases of varying degrees. Each curve corresponds to a fixed polynomial degree, ranging from 4 to 16, with points reflecting models of increasing complexity obtained through regularization. A clear trend is observed: for a given polynomial degree, increasing the number of parameters generally results in improved model accuracy (i.e., lower loss). However, diminishing returns set in, and the rate of improvement flattens. More notably, the envelope formed by the lowest loss at each level of complexity across all degrees traces an emergent Pareto front. This front captures the trade-off between model simplicity and predictive performance.
@@
Figure 6: (a) The scaling behavior of the sparse regression coefficients is the same as for coefficients of Fourier series. (b) The difference between the sparse regression fit and the real values for relative humidity of 5%.
In Fourier analysis, the magnitude of coefficients typically decays as $1/n$ (where $n$ is the order of the term) for functions of bounded variation [52] - a class that includes many naturally occurring signals and is a reasonable assumption for observational data. This decay reflects the fact that higher-order (or higher-frequency) components contribute less to the overall structure of such functions. A similar trend is observed in sparse regression using orthogonal polynomial bases, see Fig. 6(a). When coefficients are sorted by magnitude, they exhibit a clear decreasing pattern, analogous to the Fourier case, with lower-order terms capturing the dominant structure and higher-order terms refining the approximation in a controlled manner.
(a) (a)
(b) (b)

### v2

@@ Abstract
The Universal Thermal Climate Index (UTCI) is a measure of thermal comfort that quantifies how humans experience environmental conditions. Due to its robustness and versatility as a bioclimatic indicator, it has been extensively employed across a wide range of studies in bioclimatology and is increasingly used as an operational measure of outdoor thermal comfort. At the same time, calculating the UTCI value from the relevant environmental parameters is nominally not straightforward, which is why using a 6th-degree polynomial approximation has become the standard way to calculate UTCI values. At the same time, although it is computationally efficient, the error of this polynomial approximation can be substantial. The goal of this study was to develop an improved version of the polynomial approximation – one that retains comparable computational efficiency but is more robust in terms of numerical stability and substantially more accurate, particularly in reducing the frequency of larger errors. This goal was successfully achieved using sparse orthogonal regression, namely sparse regression with an orthogonal polynomial basis, which not only substantially reduces the average errors (i.e., the mean error, the mean absolute error, and the root mean square error) but also drastically reduces the frequency of large errors. By leveraging Legendre polynomial bases, approximation models could be constructed that efficiently populate a Pareto front of accuracy versus complexity and exhibit stable, hierarchical coefficient structures across varying model capacities. Training the new approximation models over only 20% of the data, with the testing performed over the remaining 80%, highlights successful generalization, with the results also being robust under bootstrapping. The decomposition effectively approximates the UTCI as a Fourier-like expansion in an orthogonal basis, yielding results near the theoretical optimum in the $L_{2}$ (least squares) sense.
@@ 1 Introduction
The Universal Thermal Climate Index (UTCI) is a measure of thermal comfort that quantifies how humans experience environmental conditions. It is derived from an advanced thermo-physiological model (Pappenberger et al., 2015) and expressed in units of temperature. The index accounts for multiple factors, including air temperature, humidity, wind speed, radiation, and clothing insulation (Bröde et al., 2012). A notable advantage of the UTCI compared to many other bioclimatic indices is its ability to represent thermal conditions in terms that are applicable to human strain under a wide range of climatic conditions (e.g., for both hot and cold conditions, Blazejczyk et al. (2012)). Based on the UTCI value, the environmental conditions can be classified into one of the ten thermal stress categories (Bröde et al., 2012), ranging from Extreme heat stress (UTCI>43UTCI>43C) to Extreme cold stress (UTCI<40UTCI<-40C).
Owing to its robustness and versatility as a bioclimatic indicator, the UTCI has been extensively employed across a wide range of studies in bioclimatology and related scientific disciplines. Its applications encompass diverse research areas, including the assessment of regional and local bioclimate characteristics, the study of urban bioclimate, recreation, tourism, and sports, epidemiological and health-related research, as well as the assessment and forecasting of bioclimatic changes (Błażejczyk and Kuchcik, 2021). The UTCI has also seen growing adoption across numerous countries as a standardized measure of outdoor thermal comfort and is increasingly integrated into routine operational meteorological forecasts. For example, within Europe, UTCI is used operationally in the Czech Republic, Italy, Poland, Portugal, and Slovenia (Di Napoli et al., 2021a; Kuzmanović et al., 2024).
At the same time, calculating the UTCI value from the relevant environmental parameters is nominally not straightforward. Namely, the UTCI is based on the Fiala multi-node model of human thermoregulation (Fiala et al., 2012). However, running the complete Fiala model is computationally expensive and requires expert knowledge to operate the complex simulation software (Bröde et al., 2012). This is the reason the authors of Bröde et al. (2012) provided two simplified approximate procedures for calculating the UTCI values that could be used in operational settings. The first approximation is based on a 4-dimensional look-up table of 104 643104\,643 accurate pre-calculated UTCI values that cover a wide range of relevant combinations of the meteorological parameters. Using this look-up table, interpolation from nearby data points can be used to determine approximate UTCI values for intermediate values of meteorological parameters. The second approximation is based on a 6th-degree regression polynomial with 210 coefficients.
Each approximation has its benefits and weaknesses. The look-up table approach is more accurate, but storing the tabulated values and searching for neighboring datapoints poses challenges to the implementation of this algorithm, while also resulting in a longer execution time compared to the other approach (Bröde, 2021a). In contrast, the polynomial approximation is less accurate, but computationally faster and substantially easier to implement in various programming languages and computational environments, as it relies on only the most common, primitive mathematical operators and does not require storing the tabulated values.
Due to its simplicity and computational efficiency, the polynomial approximation has become the standard way of calculating the UTCI values. It has been incorporated into various biclimatic software packages and libraries (e.g., the Bioklima software, Błażejczyk (2025), the Thermofeel Python library, Brimicombe et al. (2022), and the pyThermalComfort Python library Tartarini and Schiavon (2020)), as well as numerical weather prediction and reanalysis systems (e.g., the ALADIN model, Termonia et al. (2018), and the ERA5 reanalysis, Di Napoli et al. (2021b)). At the same time, the error of the polynomial approximation can be substantial. For example, when evaluated on the aforementioned look-up table of accurate UTCI values, the root-mean-square-error is about 1.1C while the frequency of absolute errors larger than 2C is about 8%, and the frequency of errors larger than 3C is about 2%. This is problematic since an error of a few degrees Celsius can increase the likelihood of misclassification of the thermal stress category, some of which span only a 6C interval.
The goal of this study is to develop an improved version of the polynomial approximation – one that has comparable computational complexity to the existing approximation but is more robust in terms of numerical stability and substantially more accurate, particularly in reducing the frequency of larger errors. To achieve this goal, symbolic and sparse regression techniques are used as tools for interpretable and efficient function approximation. We fit the UTCI offset using sparse regression on an orthogonal Legendre polynomial basis. To emphasize this key feature and distinguish it from standard sparse regression on monomials, we refer to this approach as sparse orthogonal regression.
We also note that the aim was not to derive an approximation that was as accurate as possible. For example, a sufficiently complex neural-network-based model would likely provide more accurate estimates of the UTCI values. However, such a model would also require the use of machine-learning libraries, as well as suitable Graphics Processing Units, to function efficiently. This means that its implementation in various programming languages and computational environments would be substantially more difficult. On the other hand, replacing an existing polynomial approximation with a new one is fairly straightforward, meaning that implementing the new approximation into existing biclimatic software packages/libraries and numerical weather prediction systems would be relatively easy.
2 Methods
Formally, the UTCI is defined as (Bröde et al., 2012)
$\mathit{UTCI}=\mathit{Ta}+\text{Offset}(\mathit{Ta},\mathit{va},\mathit{Tr},\mathit{rH}\text{ or }\mathit{pa}),$ (1)
where $\mathit{Ta}$ is the air temperature and the Offset is the physiologically equivalent temperature difference, representing how other environmental factors modify the effect of the thermal stress on the human body. The Offset function represents the deviation of the UTCI from the actual air temperature and depends on $\mathit{Ta}$, wind speed at 10 m ($\mathit{va}$), mean radiant temperature ($\mathit{Tr}$), which accounts for the effect of all incoming radiation, and humidity, which can be represented by either relative humidity ($\mathit{rH}$) or water vapour pressure ($\mathit{pa}$).
The dataset provided by Bröde et al. (2012) contains accurate values of the Offset function covering a wide range of environmental states. The variables and their ranges are included in Table 1. The intervals of the environmental variables also represent the domain where the sixth-degree polynomial regression approximation is considered valid (Bröde et al., 2012). Using the approximation for conditions outside of these intervals can lead to large errors and unrealistic values of the Offset function and should be avoided (Bröde, 2021a).
Variable Description Valid Normalized
name Range range
$\mathit{Ta}$ Air temperature 50-50~to+50~+50 C [1,1][-1,1]
$\mathit{va}$ Wind speed at 10 m 0.50.5~to30.3~30.3 m/s [1,1][-1,1]
$\mathit{Tr}-\mathit{Ta}$ Mean Radiant–air temperature difference 30-30~to+70~+70 C [1,1][-1,1]
$\mathit{rH}$ Relative humidity 55~to100~100 % [1,1][-1,1]
$\mathit{pa}$ Water vapour pressure 00~to5~5 kPa Not used
Table 1: Description of variables used in this study, following Bröde et al. (2012). The normalized ranges map each variable to [1,1][-1,1], with respect to the interval of validity, suitable for use with Legendre polynomial bases. We note that although $\mathit{pa}$ was not used as an input for the new approximation, its value (which can be calculated from the $\mathit{Ta}$ and $\mathit{rH}$) still needs to be located inside the specified interval of validity for the approximation to be considered valid.
@@ 2 Methods
In Fig. 1(a) we see how the UTCI Offset varies along the different environmental variables. Instead of the humidity ($\mathit{rH}$), the water vapor pressure ($\mathit{pa}$) can be used which is a nonlinear function of $\mathit{rH}$ and the air temperature ($\mathit{Ta}$). However, the variables have different distribution, see Fig. 1(b), which impacts the extent that approximations of UTCI can generalize, discussed below.
@@
Figure 1: (a) 3D plot of UTCI Offset Bröde et al. (2012) at 5% relative humidity, showing how wind speed ($\mathit{va}$), air temperature ($\mathit{Ta}$), and mean radiant temperature difference ($\mathit{Tr}-\mathit{Ta}$) combine to influence thermal stress. Color indicates the UTCI Offset magnitude across these environmental dimensions. (b) The different distributions of the water vapor pressure and relative humidity in the computed Offset dataset Bröde et al. (2012). The water vapor pressure is strongly peaked at zero, while the relative humidity is uniform across its range.
Equation discovery aims to learn interpretable mathematical expressions, either differential or algebraic equations, from measurements of the variables of a given observed system (Todorovski and Džeroski, 1997). Positioned at the intersection of symbolic machine learning and system identification, it is becoming increasingly relevant in environmental and climate science, where data-driven yet transparent models are essential (Steinmann et al., 2025; Roman, 2025b). Traditional modeling approaches rely on expert-derived formulations (Roman, 2021, 2023; Roman and Bertolotti, 2022, 2023; Roman and Palmer, 2019), but the growing complexity and volume of climate data call for automated alternatives. Symbolic regression, which iteratively combines mathematical operators and variables to fit data, forms the core of equation discovery (Bridewell et al., 2005; Todorovski and Džeroski, 2006; Džeroski et al., 2007). Most methods employ evolutionary or other (e.g., enumerative) search strategies to explore the space of candidate equations (Tanevski et al., 2016a, 2020; Mežnar et al., 2023).
Recent advances integrate probabilistic grammars to incorporate prior knowledge and constrain the search to physically meaningful expressions (Brence et al., 2020, 2023; Omejc et al., 2024). This structured approach improves both model interpretability and search efficiency, especially in domains governed by established scientific principles. Equation discovery has been applied to various environmental systems (Atanasova et al., 2006a, 2011, 2008, b), including ecosystem dynamics (Jeraj et al., 2006; Čerepnalkoski et al., 2012; Simidjievski et al., 2015, 2016; Tanevski et al., 2016b). In these settings, it can match or even surpass expert-built models while simultaneously revealing new relationships (Todorovski and Džeroski, 2001; Todorovski et al., 1998). Its ability to generate compact, interpretable, and physically plausable models makes it especially suitable for climate applications, where model transparency and adherence to physical principles are vital.
As already mentioned, the errors of the sixth-degree regression polynomial from Bröde et al. (2012) can be substantial. Fig. 2(a) shows the approximation error at 5% relative humidity, while 2(b) displays a histogram of the errors, revealing a normal distribution centered at zero, indicating minimal bias. We aim to improve upon this standard approximation using equation discovery and sparse regression methods by utilizing the accurate Offset dataset provided by Bröde et al. (2012).
@@
Figure 2: The error of the standard polynomial UTCI approximation (Bröde et al., 2012) for relative humidity of 5%. (a) The difference between the standard UTCI approximation and the accurate values of the Offset function. (b) Histogram of the differences showing a normal distribution centered at zero.
Sparse machine learning models aim to construct parsimonious predictive functions by enforcing zero-valued coefficients in high-dimensional parameter spaces, thereby performing implicit feature selection (Brunton et al., 2016). This sparsity promotes interpretability, reduces overfitting, and improves computational tractability, especially when the number of candidate predictors is large or when strong correlations exist among inputs. Sparse regression (Brunton et al., 2016), a key instantiation of this paradigm, extends linear regression with an $L_{1}$-norm regularization term—most notably in the Lasso (Least Absolute Shrinkage and Selection Operator, Reid et al., 2016) —to penalize unnecessary parameters and induce a compact representation.
@@ 3 Results and discussion
Table 2 presents a detailed comparison of model performance across a range of polynomial degrees for both standard (non-sparse) linear regression and sparse regression techniques, evaluated in the context of approximating the UTCI. The standard approximation (Bröde et al., 2012) is a sixth-degree regression polynomial model with four variables, consisting of 210 terms and achieving a root mean squared loss of $1.12^{\circ}$C. This serves as the benchmark to be matched or improved upon. It is important to note that the standard approximation does not directly employ the relative humidity ($\mathit{rH}$), but the water vapor pressure ($\mathit{pa}$), which can be derived from the relative humidity ($\mathit{rH}$) and air temperature ($\mathit{Ta}$). As we noted above, in the dataset, the relative humidity is well represented across its entire range, see Fig. 1(b), while the water vapor pressure is strongly peaked close to zero. Optimization employing the water vapor pressure ($\mathit{pa}$) as an independent variable (instead of $\mathit{rH}$) is thus poorly conditioned and leads to instability in the regression coefficients, both in simple and sparse regression. While using the pa (instead of $\mathit{rH}$) can achieve better accuracy (lower loss), it comes at the price of losing parameter consistency across optimizations with different polynomial degrees. For this reason, we report our results employing the relative humidity ($\mathit{rH}$) instead of the water vapor pressure ($\mathit{pa}$), see Table 1.
@@
Table 2: Root mean squared train loss [C], test loss [C] and the number of parameters (shown in parenthesis) in approximating the UTCI Offset. The baseline reference, labeled as “Standard,” corresponds to the sixth-degree regression polynomial model with four variables (Bröde et al., 2012). Unless otherwise stated the test loss equals the train loss. Where two loss values are reported, they indicate a notable train-test discrepancy, typically suggesting overfitting. Training is done with 20% of the data and testing is performed with 80%. Results are robust under bootstrapping.
@@
The regression methods are applied to polynomial basis expansions of increasing degree, evaluated on the basis of root mean squared test loss and number of active parameters. Unlike many studies in the literature where models are trained on the majority of the data and evaluated on a relatively small test set, our approach inverts this paradigm: training is conducted on only 20% of the available data, while performance is assessed on the remaining 80%. Despite this stringent evaluation setting, the models achieve comparable performance on both training and test sets, underscoring their strong generalization capabilities. This performance stability is further validated through bootstrapping, which reveals minimal variance in both loss metrics and selected features across resampled datasets. The reported performance metrics—such as train/test loss and number of parameters—remain stable when the model training and evaluation process is repeated on multiple random re-samplings (bootstrapped subsets) of the data. This suggests that the results are not sensitive to specific data splits and that the models generalize well across different subsets of the dataset, indicating reliability and consistency in the reported findings. These findings demonstrate the robustness and reliability of the proposed framework.
@@
Linear regression without any sparsity constraints shows improved performance at higher degrees, with test loss reducing as model capacity increases. However, this comes with a dramatic increase in the number of parameters; it reaches over 1800 coefficients by degree 12. Furthermore, the discrepancy between train and test losses at higher degrees (e.g., 0.62C vs. 0.54C at degree 12) indicates overfitting, despite the improved predictive accuracy. The resulting models are also substantially more complex, raising concerns regarding interpretation and generalization. Sparse regression with standard polynomial bases show similar performance at low degrees but fails to converge beyond the 6th degree. This indicates that enforcing sparsity in a poorly conditioned basis becomes increasingly difficult as model complexity grows.
@@
Figure 3: Loss versus number of parameters for different polynomial degrees. The regularization parameter was varied in the lasso regression to yield a Pareto front in model accuracy and complexity for each degree.
@@
In contrast, sparse regression using an orthogonal Legendre basis (or sparse orthogonal regression) exhibits superior stability and accuracy across all degrees. It outperforms the baseline 6th-degree polynomial fit from degree 8th onward, achieving a test loss of 0.88C at degree 10 with only 209 parameters—almost the same count as the original benchmark model, but with improved generalization. As the degree increases to 16, the loss reduces further to 0.60C using 424 parameters—a fraction of those used by the corresponding standard regression model. The orthogonality of the Legendre basis likely contributes to better numerical conditioning, facilitating sparse model discovery even at high degrees. These results emphasize the importance of basis selection and regularization strategy in symbolic regression tasks. Sparse methods, when combined with well-structured bases like Legendre polynomials, offer a promising path toward accurate, compact, and interpretable models in high-dimensional settings.
@@
(c) (c)
Figure 4: Parameters (or polynomial coefficients) and how they change for different polynomial degrees for (a) simple regression and (b) sparse regression (using Legendre basis). (c) Sorted sparse‐regression coefficients (Legendre basis) versus parameter index on a logarithmic x–axis show a clear, Fourier–like decay with order—approximately 1/n1/n—that is stable across model capacities (degrees 4, 8, 12, 16), indicating a hierarchical structure where lower–order terms dominate and higher–order terms provide incremental refinement.
Fig. 3 illustrates the relationship between model complexity (measured by the number of parameters) and prediction accuracy (log-scaled loss) for sparse regression models using Legendre polynomial bases of varying degrees. Each curve corresponds to a fixed polynomial degree, ranging from 4 to 16, with points reflecting models of increasing complexity obtained through regularization. A clear trend is observed: for a given polynomial degree, increasing the number of parameters generally results in improved model accuracy (i.e., lower loss). However, diminishing returns set in, and the rate of improvement flattens. More notably, the envelope formed by the lowest loss at each level of complexity across all degrees traces an emergent Pareto front (Smits and Kotanchek, 2005). This front captures the trade-off between model simplicity and predictive performance.
@@
In Fourier analysis, the magnitude of coefficients typically decays as $1/n$ (where $n$ is the order of the term) for functions of bounded variation (Stein and Shakarchi, 2011) – a class that includes many naturally occurring signals and is a reasonable assumption for observational data. This decay reflects the fact that higher-order (or higher-frequency) components contribute less to the overall structure of such functions. A similar trend is observed in sparse regression using orthogonal polynomial bases, see Fig. 4(c). When coefficients are sorted by magnitude, they exhibit a clear decreasing pattern, analogous to the Fourier case, with lower-order terms capturing the dominant structure and higher-order terms refining the approximation in a controlled manner.
This suggests that through the use of sparse regression with an orthogonal polynomial basis, we have achieved a Fourier-like decomposition of the UTCI Offset in the Legendre basis (instead of the trigonometric one). This has a number of theoretical advantages: due to the orthogonality of the basis functions, the decomposition minimizes the $L_{2}$ distance (least squares) between approximation and function, guaranteeing the best possible polynomial fit for a given model complexity (Stein and Shakarchi, 2011). Additionally, the coefficients are uncorrelated and hierarchically structured, ensuring that lower-order components remain stable as higher-order terms are added—enhancing both interpretability and numerical robustness.
Based on the analysis results and one of the initial goals (that the new approximation should have comparable computational complexity to the existing one), we selected the sparse regression model based on tenth-degree Legendre polynomials as the most suitable approximation. The final version of the new polynomial, which has 209 coefficients, was calculated using the whole dataset of tabulated values.
Fig. 5(a) shows the spatial distribution of the Offset errors for the new approximation at a fixed relative humidity of 5%. The errors are small and smoothly varying, indicating good agreement across the input space. Fig. 5(b) presents a comparison of error histograms for both the standard and new approximations. The sparse-model-based approximation produces a narrower, more sharply peaked distribution centered at zero, highlighting a reduction in error variance and suggesting better generalization. Fig. 5(c) shows the cumulative distribution of absolute errors for the two approximations. The curve for the new approximation rises more steeply and reaches higher cumulative values at lower error thresholds, indicating that a larger proportion of predictions fall within smaller error margins.
@@
(c) (c)